OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). y = 7x - 3 y = -7x + 10 y = 7x - 10 y = -1/7 x + 2

OpenStudy (anonymous):

It's already in slope - intercept form, which is y =mx + b form.

OpenStudy (anonymous):

Ah parallel :), that makes much more sense ^.^

OpenStudy (anonymous):

So for that just use the given slope (7), and use y - y1 = m(x-x1), put in the (x, y) and boom. Parallel line :). y+3 =7(x-1) y+3 = 7x -7 y = 7x-10

OpenStudy (anonymous):

Thank you